OpenStudy (anonymous):

What are the steps to finding the limit for: lim tan(6x)/sin(2x) x==>0

OpenStudy (anonymous):

have you tried l'hopital?

OpenStudy (anonymous):

No, I have no idea what that is

OpenStudy (anonymous):

i am going to guess it is 3

OpenStudy (anonymous):

rewrite with only sines and cosines, then ignore the cosine part because as \(x\to 0\) you know \(\cos(x)\to 1\)

OpenStudy (anonymous):

that leaves you with \[\lim_{x\to 0}\frac{\sin(6x)}{\sin(2x)}\] which is a trivial l'hopital problem, but if you have not got there yet (and i suspect that you have not) then you can rewrite this in the somewhat artificial and cumbersome form as \[\frac{\sin(6x)}{6x}\times \frac{2x}{\sin(2x)}\times 3\] then use the fact that \[\lim_{x\to 0}\frac{\sin(x)}{x}=1\] to get and answer of 3